Circular No. 5931

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

A0535+26 M. H. Finger, Computer Sciences Corporation; R. B. Wilson, Mar-

shall Space Flight Center, NASA; and K. S. Hagedon, Universities Space Research Association, report for the Compton Gamma Ray Observatory BATSE team: “Hard x-ray pulses in the energy range 20-120 keV are currently being detected from the transient x-ray pulsar A0535+26. The present outburst was first detected on Jan. 28. The pulsed flux remained only marginally detectable until Feb. 1. A steady increase in flux began on Feb. 3.0, when the pulsed flux doubled. The spectrum of the phaseaveraged pulsed flux on Feb. 3 is adequately fitted by an optically-thin thermal bremsstrahlung model F (E) = (A/E) exp(−E/kT), with a temperature kT = 27.7 ± 1.0 keV, and F (50 keV) = (5.1 ± 0.1) × 10^{−4} photon cm^{−2} s^{−1} keV^{−1}. The pulse shape is very broad, with a sharp dip occupying ∼ 0.2 in phase. This is the largest of a series of four consecutive outbursts of A0535+26 that has been detected by BATSE. Outbursts in 1993 March and July were reported on IAUC 5388. A weak outburst occurred in 1993 Oct.-Dec. The following preliminary binary orbital parameters have been determined from a pulse timing analysis of the previous three outbursts: P = 110.3 ± 0.3 days, T = 1993 Mar. 12.7 ± 0.6 UT, a sin i = 267 ± 13 light-s, e = 0.47 ± 0.02, and ω = 130^{o} ± 5^{o}. An intrinsic pulse period of 103.4216 ± 0.0010 s was determined for Feb. 3 using these orbital parameters.”

KR AURIGAE Visual magnitude estimates (cf. IAUC 5927): Jan. 9.75 UT, 13.8 (M.

Verdenet, Bourbon Lancy, France); 29.81, 13.4 (Verdenet); 30.78, 13.3 (Verdenet); 31.86, 13.4 (Verdenet); 31.91, 12.7 (A. Mizser, Budapest, Hungary); Feb. 6.76, 13.5 (Verdenet).

PERIODIC COMET KUSHIDA (1994a) Total visual magnitude estimates: Jan. 19.94 UT, 11.6 (K. Sárneczky,

Veresegyház, Hungary, 0.33-m reflector); 23.16, 10.1 (H. Mikuž, Ljubljana, Slovenia, 0.20-m reflector + CCD); Feb. 2.17, 10.3 (A. Hale, near Orogrande, NM, 0.20-m reflector); 3.18, 10.0 (J. E. Bortle, Stormville, NY, 0.32-m reflector); 8.18, 11.6 (C. E. Spratt, Victoria, BC, 0.20-m reflector).

1994 February 8 Daniel W. E. Green